OpenStudy (anonymous):

Use inscribed rectangles to approximate the area under g(x) = –0.5x2 + 7 for –3 ≤ x ≤ 0 and rectangle width 0.5.

OpenStudy (anonymous):

this is a Reimann sums problem \[g(x)=-0.5x^2+7\\ h=0.5\\ a=-3\qquad b=0\\ N={b-a\over h}={3\over 0.5}=6\\ \text{Let us use left-hand sums}\\ x_0=a=-3\qquad\qquad g(x_0)=-0.5(-3)^2+7=2.5\\ x_1=a+h=-2.5\quad g(x_1)=3.875\\ x_2=a+2h=-2\quad g(x_2)=5\\ \vdots\\ x_5=a+5h=-0.5\quad g(x_5)=6.875\\ \;\\ \text{approximation of the area is:}\\ A\approx h\sum_{k=0}^{N-1}g(x_k)\\ A\approx h\left[2.5+3.875+5+\ldots+6.875\right]=? \]

OpenStudy (anonymous):

you follow?

OpenStudy (anonymous):

um, no, im confused on the last part

OpenStudy (anonymous):